Here's my question: To play Wisconsin power ball, you select 5 different numbers from 1 through 59 and 1 power ball number from 1 through 39. The power ball can match a previously selected number.
So how many different ways are there to select all six numbers?
AND
What is the probability of matching all 6 numbers?

Answer by htmentor(1343)

The number of combinations of n things taken r at a time = C(n,r)
In this case we have 5 number combinations from the numbers 1 to 59:
Number of 5 number combinations = 59!/(5!*54!) = 5006386
For the powerball, there are 39 possibilities, so the total number of combinations is:
N = 5006386*39 = 195249054
So the odds in winning are 1 in 195249054